# Attention-aware Semantic Communications for Collaborative Inference

### v1

@@ I Introduction
To address the dual challenges of excessive communication overhead and limited computational resources, the concept of collaborative inference has been introduced [2, 3, 4, 6, 5, 7]. This strategy involves dividing a deep neural network (DNN) model into separate parts for the edge device and the server. Within this framework, the edge device first uses its component to extract features from the raw data and then transmits them to the server. As these extracted features are typically more compressed than the raw data, the communication cost can be aggressively reduced. The server then utilizes these features and its portion of the model to determine the final inference result, which is sent back to the device [5]. The selection of the split point is critical as it significantly impacts the computational load on the device and the communication overhead [3]. This approach is also known as split inference [8, 5] and device-edge co-inference [3, 7]. Notably, collaborative inference is closely connected to semantic communications [9, 10, 11, 12, 13, 14], considering that the extracted features are essentially semantic information tailored for the inference task.
@@
In collaborative inference scenarios, the strategy of partitioning ViT models fails to effectively reduce communication overhead. This limitation stems from the inherent architecture of ViTs, which maintains consistent layer dimensions across the entire transformer encoder [16], in contrast to convolutional neural network (CNN) models whose intermediate layer dimensions can be significantly smaller than the raw data dimensions. Consequently, partitioning a ViT model for collaborative inference still requires the transmission of significant data volumes, thereby failing to substantially mitigate the communication load.
@@
TABLE I: The Comparison of DeiT Model Parameters and Classification Accuracy on ImageNet Dataset [17]
Model Parameters (million) Classification Accuracy (%)
DeiT-Tiny 5 72.2
DeiT-Small 22 79.8
DeiT-Base 86 81.8
Our objective is to develop a collaborative inference strategy that achieves classification accuracy comparable to the server model while also minimizing communication overhead between the edge device and the server. This strategy is designed to leverage the strengths of both models: the efficiency and low resource demand of the tiny model at the edge, and the higher classification accuracy of the base model at the server. In our proposed framework, the edge device utilizes its tiny model to conduct an initial inference without transmitting the image to the server. Subsequently, the edge device assesses whether to accept this initial inference or to send the image to the server for a more accurate inference using the base model. By doing so, we aim to achieve an optimal trade-off between classification accuracy and communication cost in edge-server collaborative systems.
@@
  • Attention-aware patch selection: In cases where the edge device needs to transmit the image to the server, our strategy is to transmit only the essential patches that are crucial for classification, rather than the entire image. This selective transmission approach is guided by the attention scores from the class token [16] to the image patches, as processed by the tiny model. These attention scores indicate the relevance of each patch to the classification task. We investigate several techniques to effectively select these key patches based on the attention scores, showing that the tiny model is capable of accurately identifying the most informative patches, even in cases where it may not classify the image correctly. Our experimental results validate that this approach enables the base model to maintain its classification accuracy, although it processes only selectively transmitted patches. Hence, attention-aware patch selection can effectively reduce the communication cost between the edge device and the server while maintaining classification accuracy.
  • Entropy-aware image transmission: It is critical for the edge device to make an accurate determination between relying on the tiny model’s inference and requesting the inference of the more accurate base model. This decision is particularly important due to the communication overhead involved in transmitting images. Ideally, the edge device should transmit only those images incorrectly classified by the tiny model while avoiding the transmission of images where the initial inference is correct. To facilitate this decision, we utilize the min-entropy derived from the softmax output values of the classification head in the tiny ViT model. Our experimental findings reveal that the decisions based on the min-entropy yield higher accuracy compared to those based on the Shannon entropy.
In the context of semantic communications, our framework is aptly characterized as attention-aware semantic communications. The tiny model’s transformer encoder on the edge device acts as a semantic encoder, particularly when the decision is made to transmit the image to the server. The edge device leverages attention scores generated by the transformer encoder to identify the most essential image patches, which are critical for accurate classification. Interestingly, our findings reveal that the tiny model functions effectively as a semantic encoder, in spite of its lower classification accuracy compared to the base model.
@@ II-B Related Work
Prior work on collaborative inference has primarily focused on CNN architectures [2, 3, 4, 6, 5, 7]. The exploration of collaborative inference with transformer models has been limited because of the consistent dimension of the MSA blocks in the transformer encoder.
Recent studies have delved into on-device inference utilizing lightweight ViT models [18, 19, 20], yet achieving the accuracy of server models is challenging. To improve classification accuracy of the edge device, an ensemble strategy employing multiple edge devices was proposed in [21]. In this approach, a central edge device disseminates raw data to nearby edge devices, which then conduct inference using compact ViT models and return their intermediate inference results to the central edge device. The central edge device obtains the final inference outcome by aggregating the received intermediate results. However, this approach leads to increased communication overhead among the edge devices.
In [22], the authors proposed a collaborative inference by leveraging the masked autoencoder (MAE) [23]. The edge device selects image patches randomly and transmits them to the server. The server then reconstructs the entire image using the decoder of MAE and performs classification on this reconstructed image. In spite of the benefit of reducing computational demands on the edge devices, the classification accuracy is compromised by the random selection of image patches.
@@ III Collaborative Inference Framework Based on Transformer Models
Figure 2: The proposed collaborative inference framework utilizing pre-trained ViT models (DeiT-Tiny on client and DeiT-Base on server). (a) Case 1: The edge device (client) transmits the selected image patches to the server when the entropy of client’s inference is high. (b) Case 2: The edge device confirms its tiny model’s classification outcome when the entropy of edge device’s inference is low.
@@
Due to the consistent layer dimensions of ViTs, conventional methods of collaborative inference [2, 3, 4, 6, 5, 7], which typically partition a single CNN model, are ineffective at reducing communication costs for ViT models. As a solution, we employ a lightweight ViT model (e.g., DeiT-Tiny) at the edge device, instead of splitting a complex ViT model (e.g., DeiT-Base), as depicted in Fig. 2. The proposed inference framework establishes an efficient collaborative protocol between the edge device and the server, aiming to achieve high classification accuracy of DeiT-Base model while significantly reducing communication overhead.
In our collaborative inference framework, the edge device (client) first performs inference with its tiny model. The edge device then evaluates the entropy level of this initial inference. High entropy (low confidence) necessitates transmitting the image to the server since it indicates that the tiny model’s inference would be unreliable. In such instances, only essential patches for classification are transmitted instead of the entire image patches to minimize communication costs. The server, utilizing its complex ViT model, conducts inference based on these selected patches and sends its classification results back to the edge device, as shown in Fig. 2LABEL:sub@fig:case1. This process of selecting critical patches is governed by the proposed attention-aware patch selection rule, elaborated in Section IV.
If the initial inference’s entropy is low, the edge device confirms its classification result without further interaction with the server, as shown in Fig. 2LABEL:sub@fig:case2. Reducing reliance on the server to reduce communication costs is achieved through entropy-aware image transmission rule, detailed in Section V. By integrating these rules, our framework significantly lowers communication costs while maintaining classification accuracy comparable to the server model.
@@
2:     $y^{(i)}_{c}:=f_{c}(\mathbf{x}^{(i)})$triangleright Inference on edge device
@@
5:         $\widetilde{\mathbf{x}}^{(i)}:=\text{patch-selection}(\mathbf{x}^{(i)})$
@@
7:         ${y}^{(i)}_{s}:=f_{s}(\widetilde{\mathbf{x}}^{(i)})$triangleright Inference on server
@@
The proposed collaborative inference framework can reduce the computational complexity for the server model by limiting the inference process to only the selected patches. The computational complexity of DeiT-Base, according to [26, 27], is specified by the formula:
@@ IV Attention-aware Patch Selection
Figure 3: The visualization of attention-aware patch selection. The left is an image of ImageNet dataset that the client model (DeiT-Tiny) inaccurately classifies as ‘Hammerhead Shark’. The right shows the selected patches by attention-aware patch selection. These selected patches allow the server model (DeiT-Base) to correctly classify the image as ‘White Shark’.
@@
Figure 4: The visualization of images and attention scores. In the left column, we have the original images of ImageNet dataset. The middle column displays the attention score maps generated by DeiT-Tiny. The right column shows the patches selected by the attention-sum threshold selection. The images are labeled as ‘Brambling’, ‘Irish Setter’, and ‘Snow Leopard’, respectively.
@@ IV-A Quantifying Patch Importance
Figure 5: The comparison of image patches selected by DeiT-Tiny (middle column) and DeiT-Base (right column). The same number of patches are selected according to the mean attention scores. The left column displays the original images of ImageNet dataset. The images are labeled as ‘Miniature Schnauzer’, ‘Great Gray Owl’, and ‘Magpie’, respectively.
@@
An interesting finding is that DeiT-Tiny can act more effectively as a semantic encoder than DeiT-Base in spite of its inferior classification accuracy. Fig. 5 compares the patches selected by DeiT-Tiny (middle column) and those by DeiT-Base (right column), showing DeiT-Tiny’s superior ability to discard irrelevant image patches. This seemingly contradictory finding can be elucidated by the insights from recent work [28], which shows that large ViT models tend to allocate high attention scores to less informative background areas. It is because the large models adeptly identify patches with minimal information, such as background areas, repurposing the corresponding tokens to assimilate global image information while neglecting their spatial information. High attention scores are allocated to these repurposed patches containing global information, particularly in the background areas, as shown in Fig. 6. Although this strategy enhances the classification accuracy of larger ViT models, it compromises their effectiveness as semantic encoders. The experimental results on how model complexity affects the efficacy of semantic encoders are presented in Section VI-E.
@@
Figure 6: The comparison of attention score maps computed by DeiT-Tiny (middle column) and DeiT-Base (right column). The left column displays the original images of ImageNet dataset. The images are labeled as ‘Miniature Schnauzer’, ‘Great Gray Owl’, and ‘Magpie’, respectively.
@@ V Entropy-aware Image Transmission
where $\eta$ denotes a predetermined threshold. To access the confidence of the client’s inference, we consider two exemplary entropy measures: 1) Shannon entropy and 2) min-entropy, with their respective thresholds.
@@
Shannon entropy has been utilized in diverse applications, such as prioritizing unlabeled data inputs in active learning [34] and wireless data acquisition [35]. Our work leverages min-entropy as a metric to evaluate the quality of initial inferences made by edge devices, setting our application apart from its previous utilization.
@@ VI-A Experiment Settings
Our experiment employs the ImageNet validation dataset and resizes each image to a resolution of $224\times 224$ pixels by center cropping. An image is flattened to $N=196$ patches before the inference. We utilize the DeiT models pre-trained on the ImageNet dataset, as presented in Table I.
We evaluate the impact of varying the number of transmitted patches on communication cost and classification cost. The communication cost in our collaborative inference system is quantified by the ratio of the number of transmitted patches to the total number of patches. Consequently, if the edge device sends all image patches to the server, the communication cost is quantified as $1$.
@@
$\eta_{m}$ Expected Rate of Transmitted Images
@@
Figure 10: Comparison of the MAE-based offloading for transformer inference (MOT) [22] and our proposed collaborative inference, with the edge device employing DeiT-Tiny and the server employing DeiT-Small.

### v2

@@ I Introduction
To address the dual challenges of excessive communication overhead and limited computational resources, the concept of collaborative inference has been introduced [2, 3, 4, 5, 6]. This strategy involves dividing a deep neural network (DNN) model into separate parts for the edge device and the server. It leverages the model architecture of DNNs, where the dimensions of intermediate layers can be significantly smaller than the input dimensions. Within this framework, the edge device first uses its component to extract features from the raw data and then transmits them to the server. As these extracted features are typically more compressed than the raw data, the communication cost can be aggressively reduced. The server then utilizes these features and its portion of the model to determine the final inference result, which is sent back to the device [5]. The selection of the split point is critical as it significantly impacts the computational load on the edge device and the communication overhead [3]. This approach is also known as split inference [7, 5] and device-edge server co-inference [3, 6]. Notably, collaborative inference is closely connected to semantic communications [8, 9, 10, 11, 12, 13, 14], considering that the extracted features are essentially semantic information tailored for the inference task.
@@
In collaborative inference scenarios, the strategy of partitioning ViT models fails to effectively reduce communication overhead. This limitation stems from the inherent architecture of ViTs, which maintains consistent layer dimensions across the entire transformer encoder [16], in contrast to DNN models whose intermediate layer dimensions can be significantly smaller than the raw data dimensions. Hence, partitioning ViT models for collaborative inference cannot reduce communication overhead.
@@
TABLE I: The Comparison of DeiT Model Complexity and Classification Accuracy on the ImageNet Dataset [17]
Model Parameters Memory FLOPs Classification
(million) (MB) (G) Accuracy (%)
DeiT-Tiny 5 21.22 1.26 72.2
DeiT-Small 22 83.21 4.61 79.8
DeiT-Base 86 329.55 17.58 81.8
Our objective is to develop a collaborative inference strategy that achieves classification accuracy comparable to the server model while minimizing communication overhead between the edge device and the server. This strategy is designed to leverage the strengths of both models: the efficiency and low resource demand of the tiny model on the edge device, and the higher classification accuracy of the base model on the server. In our proposed framework, the edge device utilizes its tiny model to conduct an initial inference without transmitting the image to the server. Subsequently, the edge device assesses whether to accept this initial inference or to send the image to the server for a more accurate inference using the base model. By doing so, we aim to achieve an optimal trade-off between classification accuracy and communication cost in edge-server collaborative systems.
@@
  • Attention-aware patch selection: In cases where the edge device needs to transmit the image to the server, our strategy is to transmit only the essential patches that are crucial for classification, rather than the entire image. This selective transmission approach is guided by the attention scores from the class token to the image patches, as processed by the tiny model. These attention scores indicate the relevance of each patch to the classification task. We investigate several techniques to effectively select these important image patches using attention scores. A crucial finding is that the tiny model is capable of accurately identifying the essential patches for the inference task, even when the client classifies the image incorrectly. Our experimental results validate that this approach enables the server model to maintain its classification accuracy, although it processes only selectively transmitted patches. Hence, our attention-aware patch selection can effectively reduce the communication cost while achieving the classification accuracy of the server model.
  • Entropy-aware image transmission: It is critical for the edge device to make an accurate decision between relying on its tiny model and requesting a more accurate inference from the base model on the server. This decision significantly affects both the communication overhead and the classification accuracy. Ideally, the edge device should transmit only those images incorrectly classified by the tiny model while avoiding the transmission of images where the initial inference is correct. To facilitate this decision, we utilize the min-entropy derived from the softmax output values of the classification head in the tiny ViT model. Our experimental findings reveal that the decisions based on the min-entropy yield higher accuracy compared to those based on the Shannon entropy.
In the context of semantic communications, our framework is aptly characterized as attention-aware semantic communications. The tiny model’s transformer encoder on the edge device acts as a semantic encoder, particularly when the decision is made to transmit the image to the server. The edge device leverages attention scores generated by the transformer encoder to identify the most essential image patches, which are critical for accurate classification. Interestingly, our findings reveal that the tiny model functions effectively as a semantic encoder, in spite of its lower classification accuracy compared to the server model.
@@ II-B Related Work
Prior work on collaborative inference has primarily focused on convolutional neural network (CNN) architectures [2, 3, 4, 19, 5, 6]. The exploration of collaborative inference with transformer models has been limited because of the consistent dimension of the MSA blocks in the transformer encoder.
Recent studies have delved into on-device inference utilizing lightweight ViT models [18, 20, 21], yet achieving the accuracy of server models is challenging. To improve the classification accuracy of the edge device, an ensemble strategy employing multiple edge devices was proposed in [22]. In this approach, a central edge device disseminates raw data to nearby edge devices, which then conduct inference using compact ViT models and return their intermediate inference results to the central edge device. The central edge device obtains the final inference outcome by aggregating the received intermediate results. However, this approach leads to increased communication overhead among the edge devices.
To reduce communication overhead, recent studies have introduced collaborative inference schemes utilizing ViT models on the server, namely masked autoencoder (MAE)-based offloading for transformer inference (MOT) [23] and adaptive MOT (A-MOT) [24]. In these approaches, the edge device selects image patches randomly for transmission to the server. The server then reconstructs the entire image using the decoder of MAE and performs classification on this reconstructed image. In spite of the benefit of reducing computational demands on edge devices, the classification accuracy is compromised by the random selection of image patches.
@@
Recent studies have considered the importance of training data samples to enhance training performance in edge learning. In [27, 28], the authors attempt to identify important training data samples and allocate more communication resources to these samples for improved communication efficiency. The authors of [29] select important features of training data samples to enhance communication efficiency during the training phase. In contrast, we focus on communication-efficient collaborative inference, which specifically identifies important patches of test data samples and transmits only these essential patches. Unlike these works that aim to enhance training performance, our framework is designed to improve communication efficiency during the inference phase.
@@ III Collaborative Inference Framework Based on Transformer Models
Figure 2: The proposed collaborative inference framework utilizing pre-trained ViT models: A lightweight model (e.g., DeiT-Tiny) on the edge device and a complicated model (DeiT-Base) on the server. The edge device (client) evaluates the uncertainty of its inference results by computing the entropy. If this entropy exceeds a predetermined threshold, the edge device selects the most important patches of the image based on the attention scores and transmits these to the server.
@@
Due to the consistent layer dimensions of ViTs, conventional methods of collaborative inference [2, 3, 4, 19, 5, 6], which typically partition a single DNN model, are ineffective at reducing communication costs for ViT models. As a solution, we employ a lightweight ViT model (e.g., DeiT-Tiny) at the edge device, instead of splitting a complex ViT model (e.g., DeiT-Base), as depicted in Fig. 2. The proposed inference framework establishes an efficient collaborative protocol between the edge device and the server, aiming to achieve high classification accuracy of DeiT-Base model while significantly reducing communication overhead.
In our collaborative inference framework, the edge device (client) first performs inference with its tiny model. The edge device then evaluates the entropy level of this initial inference. High entropy (or low confidence) necessitates transmitting the image to the server since it indicates that the tiny model’s inference would be unreliable. In such instances, only essential patches for classification are transmitted instead of the entire image patches to minimize communication costs. The server, utilizing its complex ViT model, conducts inference based on these selected patches and sends its classification results back to the edge device, as shown in Fig. 2. This process of selecting critical patches is governed by the proposed attention-aware patch selection rule, elaborated in Section IV.
If the initial inference’s entropy is low, the edge device confirms its classification result without further interaction with the server, as shown in Fig. 2. Reducing reliance on the server to reduce communication costs is achieved through entropy-aware image transmission rule, detailed in Section V. By integrating these rules, our framework significantly lowers communication costs while maintaining classification accuracy comparable to the server model.
@@
2:     $y^{(i)}_{c}\leftarrow f_{c}(\mathbf{x}^{(i)})$triangleright Inference on edge device
@@
5:         $\widetilde{\mathbf{x}}^{(i)}\leftarrow\text{patch-selection}(\mathbf{x}^{(i)})$
@@
7:         ${y}^{(i)}_{s}\leftarrow f_{s}(\widetilde{\mathbf{x}}^{(i)})$triangleright Inference on server
@@
The proposed collaborative inference framework can reduce the computational complexity for the server model by limiting the inference process to only the selected patches. The computational complexity of DeiT-Base, according to [30, 31], is given by
@@ IV Attention-aware Patch Selection
Figure 3: The visualization of attention-aware patch selection. The left is an image of the ImageNet dataset that the client model (DeiT-Tiny) inaccurately classifies as ‘Hammerhead Shark’. The right shows the selected patches by attention-aware patch selection. These selected patches allow the server model (DeiT-Base) to correctly classify the image as ‘White Shark’.
@@
Figure 4: The visualization of images and attention scores. In the left column, we have the original images of the ImageNet dataset. The middle column displays the attention score maps generated by DeiT-Tiny. The right column shows the patches selected by the attention-sum threshold selection. The images are labeled as ‘Brambling’, ‘Irish Setter’, and ‘Snow Leopard’, respectively.
@@ IV-A Quantifying Patch Importance
Figure 5: The comparison of image patches selected by DeiT-Tiny (middle column) and DeiT-Base (right column). The same number of patches are selected according to the mean attention scores. The left column displays the original images of the ImageNet dataset. The images are labeled as ‘Miniature Schnauzer’, ‘Great Gray Owl’, and ‘Magpie’, respectively.
@@
An interesting finding is that DeiT-Tiny can act more effectively as a semantic encoder than DeiT-Base in spite of its inferior classification accuracy. Fig. 5 compares the patches selected by DeiT-Tiny (middle column) and those by DeiT-Base (right column), showing DeiT-Tiny’s superior ability to discard irrelevant image patches. This seemingly contradictory finding can be elucidated by the insights from recent work [32], which shows that large ViT models tend to allocate high attention scores to less informative background areas. It is because the large models adeptly identify patches with minimal information, such as background areas, repurposing the corresponding patches to assimilate global image information while neglecting their spatial information. High attention scores are allocated to these repurposed patches containing global information, particularly in the background areas, as shown in Fig. 6. Although this strategy enhances the classification accuracy of larger ViT models, it compromises their effectiveness as semantic encoders. The experimental results on how model complexity affects the efficacy of semantic encoders are presented in Section VI-E.
@@
Figure 6: The comparison of attention score maps computed by DeiT-Tiny (middle column) and DeiT-Base (right column). The left column displays the original images of the ImageNet dataset. The images are labeled as ‘Miniature Schnauzer’, ‘Great Gray Owl’, and ‘Magpie’, respectively.
@@ V Entropy-aware Image Transmission
where $\eta$ denotes a predetermined threshold.
To assess the confidence of the client’s inference, we consider two exemplary entropy measures: 1) Shannon entropy and 2) min-entropy, with their respective thresholds.
@@
The entropy has been utilized in diverse applications, such as prioritizing unlabeled data inputs in active learning [38] and optimizing wireless data acquisition for edge learning [28]. These works typically enhance training procedures using the entropy values calculated by complicated server models. In contrast, our approach utilizes the min-entropy to assess the uncertainty of initial inferences made by a tiny model on the edge device instead of a complex server model. This use of uncertainty metrics aims to minimize unnecessary image transmissions, thereby reducing communication overhead.
@@ VI-A Experiment Settings
Our experiments employ the ImageNet validation dataset and resize each image to a resolution of $224\times 224$ pixels by center cropping. An image is flattened to $N=196$ patches before the inference.
We deploy DeiT-Tiny on the edge device and DeiT-Base on the server since resource-constrained edge devices have challenges in employing complicated models such as DeiT-Base. The model complexity of these models is compared in Table I. DeiT-Tiny can be viable for deployment on edge devices such as NVIDIA Jetson Nano [22], Raspberry Pi 4B [39], and iPhone 12 [18]. In contrast, DeiT-Base is generally considered inappropriate due to its substantial memory consumption and prolonged inference latency. For instance, the authors of [22] explicitly state that DeiT-Base is inadequate for deployment on NVIDIA Jetson Nano due to its substantial memory consumption and computational complexity. Similarly, the authors of [18] emphasize memory consumption as a critical factor for edge device deployment, considering only lightweight models such as DeiT-Tiny and PiT [40] with approximately 3–6 million parameters.
We evaluate the impact of varying the number of transmitted patches on communication cost and classification accuracy. The communication cost in our collaborative inference system is quantified by the ratio of the number of transmitted patches to the total number of patches. Consequently, if the edge device sends all image patches to the server, the communication cost is quantified as $1$.
@@
By controlling the threshold values of $\eta$ and $\delta$, our proposed collaborative inference framework can effectively manage communication resources. In scenarios with reliable communication channels, lower values of $\eta$ and $\delta$ can be selected to increase data transmissions and maximize classification accuracy. When the channel quality is worse and communication resources are limited, higher values of $\eta$ and $\delta$ can be set to balance classification accuracy with the available communication resource budget. This strategy offers flexibility in adapting to varying communication channel conditions, thereby enhancing the efficiency of communication resource management.
@@
$\eta_{m}$ Expected Ratio of Transmitted Images
@@
Figure 10: Comparison of previous methods and our collaborative inference framework, utilizing DeiT-Tiny on the edge device and DeiT-Small on the server. The experiment is conducted on the ImageNet dataset.
@@
Fig. 10 compares our collaborative inference framework with several existing methods, including server-driven transmission (SDT) [41], image super-resolution (ISR) [42], masked autoencoder (MAE)-based offloading for transformer inference (MOT) [23], and adaptive MOT (A-MOT) [24]. SDT transmits low-quality images initially and then offloads high-quality content from the target area based on the server’s feedback [41]. ISR transmits a low-quality image and then reconstructs a high-resolution image by super-resolution method [42]. The server model performs inference on this reconstructed high-resolution image. In MOT and A-MOT, the edge device randomly selects image patches for transmission to the server. The server then reconstructs the entire image using the decoder of MAE [43] and performs classification on this reconstructed image. The experimental results show that our proposed collaborative inference framework significantly improves performance by employing the tiny model on the edge device, which adeptly transmits essential image patches for classification.